Find the modulus and the principal argument of each of the given complex numbers. (a) 3 − 4i, (b) −2 + i, (c) 1/1 + i √3 (d) (7 − i)/(-4 - 3i) (e) 5(cos π/3 + isin π/3) (f) cos 2π/3 − sin 2π/3.
1
Expert's answer
2020-05-28T18:17:52-0400

The formula for modulus (r) and principal argument ("\theta") of a complex number (z = x + iy) is :


r = "\mid"z"\mid" = "\sqrt{\smash[b]{x² + y²}}"

"\theta" = arctan(y/x)


(a) z = 3 - 4i

Modulus, r = "\sqrt{\smash[b]{3² + (-4)²}}"

r = "\sqrt{\smash[b]{9 + 16}}" = "\sqrt{\smash[b]{25}}"

"\therefore" r = 5


Principal argument, "\theta" = arctan(-4/3)

"\therefore" "\theta" = -53.13"\degree"

(b) z = -2 + i

Modulus, r = "\sqrt{\smash[b]{(-2)² + (1)²}}"

r = "\sqrt{\smash[b]{4 + 1}}"

"\therefore" r = "\sqrt{\smash[b]{5}}"


Using arctan(1/(-2))"= -26.56\degree"

Since the given complex number lies in the second quadrant,

Therefore, the principal argument is "\theta = (- 26.56 + 180)\degree"

"\therefore \theta = 153.44\degree"


(c) z = 1/(1 + i"\sqrt{\smash[b]{3}}" )

"z = (1 - i\sqrt{\smash[b]{3}})/((1 + i\sqrt{\smash[b]{3}})(1 - i\sqrt{\smash[b]{3}}))"

"z = (1 - i\sqrt{\smash[b]{3}})/(1² - (i\sqrt{\smash[b]{3}})²)"

"z = (1 - i\sqrt{\smash[b]{3}})/(1 + 9)"

"z = (1 - i\sqrt{\smash[b]{3}})/10"

"z = (1/10) - i(\sqrt{\smash[b]{3}}/10)"


Modulus, r = "\sqrt{\smash[b]{(1/10)² + (\sqrt{\smash[b]{3}}/10)²}}"

r = "\sqrt{\smash[b]{(1/100) + (3/100)}}"

r = "\sqrt{\smash[b]{(4/100)}}"

"\therefore" r = (2/10)


Principal argument, θ = arctan"( -(\sqrt{\smash[b]{3}}/10)/(1/10)"

"\theta = arctan(- \sqrt{\smash[b]{3}})"

"\theta = - ( \pi / 3)"

"\therefore" "\theta = - 60\degree"


(d) z = (7 − i)/(-4 - 3i)

z = ((7 - i)(-4 + 3i))/((-4 - 3i)(-4 + 3i))

z = (-28 + 21i +4i + 3)/((-4)² - (3i)²)

z = (-25 + 25i)/(16 + 9)

z = (-25 + 25i)/25

z = -1 + i


Modulus, r = "\sqrt{\smash[b]{(-1)² + 1²}}"

r = "\sqrt{\smash[b]{1 + 1}}"

"\therefore r= \sqrt{\smash[b]{2}}"


Using arctan(1/(-1)) = arctan(-1)= -45"\degree" .

Since the given complex number lies in the second quadrant,

Therefore, the principal argument is "\theta = (-45 + 180)°"

"\therefore\theta= 135°"


(e) z = "5(cos(\pi/3) +isin(\pi/3))"

z= "5cos(\pi/3) +i5sin(\pi/3)"


Modulus, r = "\sqrt{\smash[b]{(5cos(π/3))² + (5sin(π/3))²}}"

"r = \sqrt{\smash[b]{5²(cos²(π/3) + sin²(π/3))}}"

"r = 5 \sqrt{\smash[b]{1}}"

"\therefore" r = 5


Principal argument, θ = arctan(5sin(π/3)/5cos(π/3))

θ = arctan(tan(π/3))

θ = "\pi/3"

"\therefore\theta" = 60"\degree"


(f) z = cos 2π/3 − sin 2π/3

z = (-1/2) - ("\sqrt{\smash[b]{3}}/2" )

"z = - (1 + \sqrt{\smash[b]{3}})/2"


Modulus, r = "|-(1+\sqrt{3})/2|=(1+\sqrt{3})/2"

"\therefore" "r =(1+\sqrt{3})/2"


Since the given complex number is negative real number,

Therefore, the principal argument is "\theta=180°"

"\therefore\theta= 180°"